# Ferromagnetism Mechanism in a Geometrically Frustrated Triangular Lattice

### v1

@@ III Itinerant Ferromagnetism at Electron Doping $\delta=1/2$
Fig. S5: Itinerant ferromagnetism signature for the $t_{\mathrm{QQ}}-U$ model. (a) The static spin structure factor $S(\mathbf{q})$. (b) The electron momentum distribution $n(\mathbf{k})$. The black dots represent the accessible momenta in the Brillouin zone (white lines). Interpolation has been applied in the contour plot. Here, $N=8\times 4$, $U/t=10$, $\delta=1/2$.
@@
Fig. S7: Contour plots of the static spin structure factor $S(\mathbf{q})$ for the Hubbard model, illustrating two distinct iSDW phases. Panels (a) and (c) show phases at smaller doping, while panels (b) and (d) depict phases at larger doping.
IV Itinerant Ferromagnetism of $t_{\mathrm{QQ}}-U$ model
@@
Here, the projection operator is $\hat{Q}_{\mathbf{i},{\sigma}}\equiv n_{\mathbf{i},{\sigma}}$ and Eq. (S12) describes the NN hoppings between a doubly occupied site and a singly occupied site with hopping amplitude $t_{\mathrm{QQ}}$ [see Fig. 2(c) in the main text], i.e., the doublon-singlon exchange. In Fig. S5, we examine both the static spin structure factor $S(\mathbf{q})$ and electron momentum distribution $n(\mathbf{k})$ at electron doping $\delta=1/2$ for $U/t=10$, where the pronounced peak at $\Gamma$ and the well-defined Fermi surface are consistent with the itinerant ferromagnetism observed in the Hubbard model at the same doping and the same coupling strength $U/t=10$.
V Evolution of Magnetism with Doping for $L_{y}=4$
@@
For the fully polarized phase of the Hubbard model with $U/t=10$, $\delta\approx 1/2$, $L_{y}=4$, the average spin correlations $\bar{S}(r=1)$ [yellow triangles in Fig. S6(a)] align well with those [blue circles] of the $t_{\mathrm{QQ}}-U$ model Eq. (S12). This alignment underscores a deep connection between the ferromagnetic mechanisms inherent to both cases, and thus highlights the significant influence of doublon-singlon exchange in inducing ferromagnetism in the Hubbard model with finite $U/t$. In addition, for light electron doping around $\delta\approx 1/16$, the $\bar{S}(r=1)$ of the $t_{\mathrm{QQ}}-U$ model [blue circles in Fig. S6(a)] closely matches with the fully polarized value (dashed black line) $m_{s,\mathrm{F}}^{2}(\mathbf{\Gamma})\equiv[(1-\delta)/2]^{2}$, yet deviation becomes apparent around a doping range of $43/16\lesssim\delta\lesssim 1/$ 4, the same as the ferromagnetic feature of the Hubbard model at $U/t=\infty$ for $L_{y}=4$, as shown by the Fig. S6(b).
@@
Figure S6(b) presents the magnetic phase diagram as a function of $U/t$ and electron doping $\delta$ for $L_{y}=4$, identifying two distinct ferromagnetic phases at $U=\infty$ around doping $\delta\approx 1/16$ and $\delta\approx 1/2$, separated by a doping range of $43/16\lesssim\delta\lesssim 1/$ 4. The pattern of $S(\mathbf{q})$ at such separation [see Fig. S7(a)] is similar to that with finite $U$ [see Fig. S7(c)], indicating that the iSDW at small doping extends from finite $U$ to $U=\infty$ smoothly for the case of $L_{y}=4$. This numerical observation suggests the subtle competition between the iSDW and FM at intermediate doping on narrower cylinders like $L_{y}=4$ and the final stabilization of FM on wider cylinders like $L_{y}=6$ [see Fig. 4(b) in the main text]. The notable separations of ferromagnetic phases around $\delta\approx 1/16$, $\delta\approx 1/2$ and $\delta\to 1$ for a finite system with $L_{y}=4$ hint at diverse underlying mechanisms, potentially differentiating the observed ferromagnetism around $\delta\approx 1/2$ from the Nagaoka ferromagnetism with $\delta\to 0$ and the Mielke’s flat-band ferromagnetism or Müller-Hartmann ferromagnetism with $\delta\to 1$ for a finite size system, even though such separations of the ferromagnetic phases are absent for $L_{y}=6$ [see Fig. 4(b) in the main text] at least for the largest length $L_{x}=16$ that we have simulated. On the other hand, as the effective interaction $U/t$ decreases from $\infty$ to $U/t=10$ for $L_{y}=4$, the ferromagnetic region around lighter ($\delta\approx 1/16$) and intermediate ($\delta\approx 1/2$) doping disappears and shrinks [Fig. S6(b)], respectively, yielding a ferromagnetic landscape closely aligned with the $U/t=10$ case as depicted in Fig. S6(a). Additionally, for larger doping $15/8\lesssim\delta<$ 1, a distinct pattern in $S(\mathbf{q})$ indicates the presence of another iSDW phase [see Fig. S7(b,d)].
@@
We further present the $S(\mathbf{q})$ outside the ferromagnetism phase at small doping with finite $U$ and larger doping $15/8\lesssim\delta<$ 1 in Fig. S7. Our examination reveals two distinct features in $S(\mathbf{q})$, each indicative of a different iSDW phase.

### v2

@@ II Perturbation and the wave function analysis for ferromagnetic mechanism
Fig. S5: Itinerant ferromagnetism signature for the $\boldsymbol{t_{\mathrm{QQ}}-U}$ model. (a) The static spin structure factor $S(\mathbf{q})$. (b) The electron momentum distribution $n(\mathbf{k})$. The black dots represent the accessible momenta in the Brillouin zone (white lines). Interpolation has been applied in the contour plot. Here, $N=8\times 4$, $U/t=10$, $\delta=1/2$.
We also offer an alternative wave function analysis in the large $U$ limit of the Hubbard model with $t>0$. We focus on $S_{z}=0$ sector because the ground state resides in this sector regardless of whether the system exhibits ferromagnetism. The relevant basis states and the transitions between them are depicted in Fig. S3. In Fig. S3, we carefully account for the signs of the hopping matrix element by considering the construction of the basis, i.e., the requirement of the normal order for the creation operators acting on the vacuum state. For example:
$|\uparrow,\downarrow,\uparrow\downarrow\rangle=c_{1,\uparrow}^{\dagger}c_{2,\downarrow}^{\dagger}c_{3,\uparrow}^{\dagger}c_{3,\downarrow}^{\dagger}|0\rangle.$ (S1)
To determine the ground state $\left|\Psi_{0}^{e}\right\rangle$ for the electron-doped case, we impose the system’s symmetry and minimize the energy by choosing appropriate superpositions of basis states. The signs of the coefficients are determined based on the following principle: If the hopping matrix element connecting two configurations is negative (positive), superposing these configurations with the same (opposite) sign leads to the minimization of the energy. With these considerations, therefore, the ground state must have the following form
$|\Psi_{0}^{e}\rangle\propto$ $\frac{\sqrt{2}}{2}\left(|\uparrow,\downarrow,\uparrow\downarrow\rangle+|\downarrow,\uparrow,\uparrow\downarrow\rangle\right)$ (S2)
$-\frac{\sqrt{2}}{2}\left(|\uparrow,\uparrow\downarrow,\downarrow\rangle+|\downarrow,\uparrow\downarrow,\uparrow\rangle\right)$
$+\frac{\sqrt{2}}{2}\left(|\uparrow\downarrow,\uparrow,\downarrow\rangle+|\uparrow\downarrow,\downarrow,\uparrow\rangle\right).$
The state $\left|\Psi_{0}^{e}\right\rangle$ is symmetric under the exchange of spins and sites, thus corresponds to a triplet state, indicating ferromagnetic ordering. Conversely, for the hole-doped case, the transitions between the basis states are illustrated in Fig. S4. Following the same methodology, the ground state $|\Psi_{0}^{h}\rangle$ is constructed as:
$|\Psi_{0}^{h}\rangle\propto$ $\frac{\sqrt{2}}{2}\left(|\uparrow,\downarrow,0\rangle-|\downarrow,\uparrow,0\rangle\right)$ (S3)
$+\frac{\sqrt{2}}{2}\left(|\uparrow,0,\downarrow\rangle-|\downarrow,0,\uparrow\rangle\right)$
$+\frac{\sqrt{2}}{2}\left(|0,\uparrow,\downarrow\rangle-|0,\downarrow,\uparrow\rangle\right).$
In this case, the coefficients are chosen to produce antisymmetric combinations under the exchange of spins, resulting in a singlet state with total spin $S=0$.
@@
III Itinerant Ferromagnetism of $t_{\mathrm{QQ}}-U$ model
@@
Here, the projection operator is $\hat{Q}_{\mathbf{i},{\sigma}}\equiv n_{\mathbf{i},{\sigma}}$ and Eq. (S4) describes the NN hoppings between a doubly occupied site and a singly occupied site with hopping amplitude $t_{\mathrm{QQ}}$ [see Fig. 3(c) in the main text], i.e., the doublon-singlon exchange. In Fig. S5, we examine both the static spin structure factor $S(\mathbf{q})$ and electron momentum distribution $n(\mathbf{k})$ at electron doping $\delta=1/2$ for $U/t=10$, where the pronounced peak at $\boldsymbol{\Gamma}$ and the well-defined Fermi surface are consistent with the itinerant ferromagnetism observed in the Hubbard model at the same doping and the same coupling strength $U/t=10$.
IV Evolution of Magnetism with Doping
@@
Fig. S7: Contour plots of the static spin structure factor $\boldsymbol{S(\mathbf{q})}$ for the Hubbard model, illustrating two distinct iSDW phases. Panels (a) and (c) show phases at smaller doping, while panels (b) and (d) depict phases at larger doping.
@@
For the fully polarized phase of the Hubbard model with $U/t=10$, $\delta\approx 1/2$, $L_{y}=4$, the average spin correlations $\bar{S}(r=1)$ [yellow triangles in Fig. S6(a)] align well with those [blue circles] of the $t_{\mathrm{QQ}}-U$ model Eq. (S4). This alignment underscores a deep connection between the ferromagnetic mechanisms inherent to both cases, and thus highlights the significant influence of doublon-singlon exchange in inducing ferromagnetism in the Hubbard model with finite $U/t$. In addition, for light electron doping around $\delta\approx 1/16$, the $\bar{S}(r=1)$ of the $t_{\mathrm{QQ}}-U$ model [blue circles in Fig. S6(a)] closely matches with the fully polarized value (dashed black line) $m_{s,\mathrm{F}}^{2}(\mathbf{\Gamma})\equiv[(1-\delta)/2]^{2}$, yet deviation becomes apparent around a doping range of $3/16\lesssim\delta\lesssim 1/4$, the same as the ferromagnetic feature of the Hubbard model at $U/t=\infty$ for $L_{y}=4$, as shown by the Fig. S6(b).
@@
Figure S6(b) presents the magnetic phase diagram as a function of $U/t$ and electron doping $\delta$ for $L_{y}=4$, identifying two distinct ferromagnetic phases at $U=\infty$ around doping $\delta\approx 1/16$ and $\delta\approx 1/2$, separated by a doping range of $3/16\lesssim\delta\lesssim 1/4$. The pattern of $S(\mathbf{q})$ at such separation [see Fig. S7(a)] is similar to that with finite $U$ [see Fig. S7(c)], indicating that the iSDW at small doping extends from finite $U$ to $U=\infty$ smoothly for the case of $L_{y}=4$. This numerical observation suggests the subtle competition between the iSDW and FM at intermediate doping on narrower cylinders like $L_{y}=4$ and the final stabilization of FM on wider cylinders like $L_{y}=6$ [see Fig. 5(a) in the main text]. The notable separations of ferromagnetic phases around $\delta\approx 1/16$, $\delta\approx 1/2$ and $\delta\to 1$ for a finite system with $L_{y}=4$ hint at diverse underlying mechanisms, potentially differentiating the observed ferromagnetism around $\delta\approx 1/2$ from the Nagaoka ferromagnetism with $\delta\to 0$ and the Mielke’s flat-band ferromagnetism or Müller-Hartmann ferromagnetism with $\delta\to 1$ for a finite size system, even though such separations of the ferromagnetic phases are absent for $L_{y}=6$ [see Fig. 5(a) in the main text] at least for the largest length $L_{x}=16$ that we have simulated. On the other hand, as the effective interaction $U/t$ decreases from $\infty$ to $U/t=10$ for $L_{y}=4$, the ferromagnetic region around lighter ($\delta\approx 1/16$) and intermediate ($\delta\approx 1/2$) doping disappears and shrinks [Fig. S6(b)], respectively, yielding a ferromagnetic landscape closely aligned with the $U/t=10$ case as depicted in Fig. S6(a). Additionally, for larger doping $5/8\lesssim\delta<1$, a distinct pattern in $S(\mathbf{q})$ indicates the presence of another iSDW phase [see Figs. S7(b) and (d)].
@@
We further present the $S(\mathbf{q})$ outside the ferromagnetism phase at small doping with finite $U$ and larger doping $5/8\lesssim\delta<1$ in Fig. S7. Our examination reveals two distinct features in $S(\mathbf{q})$, each indicative of a different iSDW phase.
V RPA analysis for Stabilization of Ferromagnetism via Magnetic Anisotropy
In this section, we create a spin-wave excitation gap by introducing magnetic anisotropy into the system, which stabilize the FM. Such stabilization mechanisms are relevant not only for fundamental physics but also for the development of spintronic devices based on high-temperature ferromagnetic materials, as reported in recent experiments Gong et al. (2017); Bonilla et al. (2018); Deng et al. (2018); Gibertini et al. (2019); Gong and Zhang (2019); Bedoya-Pinto et al. (2021); Wang et al. (2022). As an example, we consider the following anisotropic term added to the Hubbard Hamiltonian
$\delta H=-J_{3}\sum_{\langle\mathbf{i}\mathbf{j}\rangle}S_{\mathbf{i}}^{z}S_{\text{$\mathbf{j}$}}^{z},$ (S5)
which can be readily realized by experiments, such as in a multilayer system. Here we require $J_{3}>0$. The anisotropic term in Eq. (S5) breaks the spin $SU(2)$ symmetry. In the following, we apply the random phase approximation (RPA) on the HF ground state to show that the anisotropic term (S5) can induce a finite gap of the spin-wave spectrum, which then protects the ferromagnetic ground state at finite temperatures. The gap can be expressed as
$\Delta=2zJ_{3},$ (S6)
where $z$ is the coordinate number.
V.1 Spin-wave dispersion without magnetic anisotropy
We illustrate the instability of FM phase based on the Hartree-Fock (HF) approximation. We start with the HF Hamiltonian for a FM state with magnetism $m$
$H_{\mathrm{HF}}$ $=\sum_{\mathbf{k},\sigma}(\epsilon_{\mathbf{k}}-\mu)c_{\mathbf{k},\sigma}^{\dagger}c_{\mathbf{k},\sigma}-\frac{1}{2}Um\sum_{\mathbf{k}}c_{\mathbf{k},\alpha}^{\dagger}{\sigma_{\alpha\beta}^{z}}c_{\mathbf{k},\beta}$
$=\sum_{\mathbf{k},\sigma}(\epsilon_{\mathbf{k}}-\mu-\frac{1}{2}Um\sigma)c_{\mathbf{k},\sigma}^{\dagger}c_{\mathbf{k},\sigma}\equiv\sum_{\mathbf{k}\sigma}\xi_{\mathbf{k},\sigma}c_{\mathbf{k},\sigma}^{\dagger}c_{\mathbf{k},\sigma},$
where $\sigma=(\sigma^{x},\sigma^{y},\sigma^{z})$ is the Pauli matrix. We denote the FM ground state of the HF Hamiltonian as $|\Omega\rangle$. We can uncover the spin wave excitation by studying the charge and spin susceptibility
$\chi^{00}(\mathbf{q},t)$ $=\frac{i}{2N}\langle T\rho_{\mathbf{q}}(t)\rho_{-\mathbf{q}}(0)\rangle,$ (S7)
$\chi^{ij}(\mathbf{q},t)$ $=\frac{i}{2N}\langle TS_{\mathbf{q}}^{i}(t)S_{-\mathbf{q}}^{j}(0)\rangle,$ (S8)
with the density and spin operators $\rho_{\mathbf{q}}$ and $S_{\mathbf{q}}^{i}$ ($i=x,y,z$)
$\rho_{\mathbf{q}}=\sum_{\mathbf{k},\sigma}c_{\mathbf{k}+\mathbf{q},\sigma}^{\dagger}c_{\mathbf{k},\sigma},\quad S_{\mathbf{q}}^{i}=\sum_{\mathbf{k},\alpha,\beta}c_{\mathbf{k}+\mathbf{q},\alpha}^{\dagger}c_{\mathbf{k},\beta}\sigma_{\alpha\beta}^{i}.$ (S9)
For the notation simplicity, we ignore an extra $\frac{1}{2}$ factor in $S_{\mathbf{q}}^{i}$.
Regarding the HF Hamiltonian, we can first define the charge and spin susceptibility
$\chi_{0}^{00}(\mathbf{q},\omega)$ $=\int_{0}^{\infty}e^{-i\omega t}\frac{i}{2N}\langle\Omega|T\rho_{\mathbf{q}}(t)\rho_{-\mathbf{q}}(0)|\Omega\rangle dt,$ (S10)
$\chi^{ij}_{0}(\mathbf{q},\omega)$ $=\int_{0}^{\infty}e^{-i\omega t}\frac{i}{2N}\langle\Omega|TS^{i}_{\mathbf{q}}(t)S^{j}_{-\mathbf{q}}(0)|\Omega\rangle dt.$ (S11)
In details, we can conduct the derivations
$\chi_{0}^{00}(\mathbf{q},\omega)$ $=\int_{0}^{\infty}dte^{-i\omega t}\frac{i}{2N}\langle\Omega|T\rho_{\mathbf{q}}(t)\rho_{-\mathbf{q}}(0)|\Omega\rangle$
$=-\frac{1}{2N}\sum_{\mathbf{k}}\sum_{\sigma}\frac{n_{\mathbf{k}+\mathbf{q},\sigma}-n_{\mathbf{k},\sigma}}{\omega+\xi_{\mathbf{k},\sigma}-\xi_{\mathbf{k}+\mathbf{q},\sigma}},$ (S12)
and
$\small\begin{split}&\chi_{0}^{ij}(\mathbf{q},\omega)\\ =&\sum_{\mathbf{k}\mathbf{k}^{\prime}}\sum_{\alpha\beta}\sum_{\alpha^{\prime}\beta^{\prime}}\langle\Omega|Tc_{\mathbf{k}+\mathbf{q},\alpha}^{\dagger}c_{\mathbf{k},\beta}\sigma_{\alpha\beta}^{i}c_{\mathbf{k}^{\prime}-\mathbf{q},\alpha^{\prime}}^{\dagger}c_{\mathbf{k}^{\prime},\beta^{\prime}}\sigma_{\alpha^{\prime}\beta^{\prime}}^{j}|\Omega\rangle\\ =&-\sum_{\mathbf{k}\mathbf{k}^{\prime}}\sum_{\alpha\beta}\sum_{\alpha^{\prime}\beta^{\prime}}\sigma_{\alpha\beta}^{i}\sigma_{\alpha^{\prime}\beta^{\prime}}^{j}\langle\Omega|c_{\mathbf{k}+\mathbf{q},\alpha}^{\dagger}c_{\mathbf{k}^{\prime},\beta^{\prime}}|\Omega\rangle\langle\Omega|c_{\mathbf{k}^{\prime}-\mathbf{q},\alpha^{\prime}}^{\dagger}c_{\mathbf{k},\beta}|\Omega\rangle\\ =&-\sum_{\mathbf{k}}\sum_{\alpha\beta}\sum_{\alpha^{\prime}\beta^{\prime}}\sigma_{\alpha\beta}^{i}\sigma_{\alpha^{\prime}\beta^{\prime}}^{j}\delta_{\alpha\beta^{\prime}}\delta_{\alpha^{\prime}\beta}\frac{n_{\mathrm{FD}}(\xi_{\mathbf{k}+\mathbf{q},\alpha})-n_{\mathrm{FD}}(\xi_{\mathbf{k},\beta})}{\omega+\xi_{\mathbf{k},\alpha}-\xi_{\mathbf{k}+\mathbf{q},\beta}}\\ =&-\delta^{ij}\sum_{\mathbf{k}}\sum_{\sigma}\frac{n_{\mathrm{FD}}(\xi_{\mathbf{k},\sigma})-n_{\mathrm{FD}}(\xi_{\mathbf{k}+\mathbf{q},\bar{\sigma}})}{\omega+\xi_{\mathbf{k},\sigma}-\xi_{\mathbf{k}+\mathbf{q},\bar{\sigma}}}\\ =&\begin{cases}\chi_{0}^{00}(\mathbf{q},\omega)&i=j=z\\ \delta_{ij}\chi_{0}^{+-}(\mathbf{q},\omega)&i,j=x,y\end{cases}\end{split}$}
where $n_{\mathrm{FD}}(\xi)=1/(e^{\beta\xi}+1)$ is the Fermi-Dirac distribution. To further obtain the spin wave, we can utilize RPA, which leads the susceptibility to be
$\chi^{00}_{\mathrm{RPA}}(\mathbf{q},\omega)$ $=\frac{\chi_{0}^{00}(\mathbf{q},\omega)}{1-U\chi_{0}^{00}(\mathbf{q},\omega)},$ (S13)
$\chi^{xx,yy}_{\mathrm{RPA}}(\mathbf{q},\omega)$ $=\frac{\chi_{0}^{+-}(\mathbf{q},\omega)}{1-U\chi_{0}^{+-}(\mathbf{q},\omega)}~.$ (S14)
with the denominators
$1-U\chi_{0}^{00}(\mathbf{q},\omega)$
$=$ $1+\frac{U}{2N}\sum_{\mathbf{k}}\sum_{\sigma}\frac{n_{\mathrm{FD}}(\xi_{\mathbf{k},\sigma})-n_{\mathrm{FD}}(\xi_{\mathbf{k}+\mathbf{q},\sigma})}{\omega+\xi_{\mathbf{k},\sigma}-\xi_{\mathbf{k}+\mathbf{q},\sigma}}$ (S15)
$1-U\chi_{0}^{+-}(\mathbf{q},\omega)$
$=$ $1+\frac{U}{2N}\sum_{\mathbf{k}}\sum_{\sigma}\frac{n_{\mathrm{FD}}(\xi_{\mathbf{k},\sigma})-n_{\mathrm{FD}}(\xi_{\mathbf{k}+\mathbf{q},\bar{\sigma}})}{\omega+\xi_{\mathbf{k},\sigma}-\xi_{\mathbf{k}+\mathbf{q},\bar{\sigma}}}.$ (S16)
In particular, at $q=0$, the conditions where the two denominators vanish recover the self-consistent equation for the Hartree-Fock parameters,
$\frac{1}{U}$ $=-\frac{1}{2}\sum_{\mathbf{k}}\sum_{\sigma}\lim_{q\rightarrow 0}\frac{n_{\mathrm{FD}}(\xi_{\mathbf{k},\sigma})-n_{\mathrm{FD}}(\xi_{\mathbf{k}+\mathbf{q},\sigma})}{\xi_{\mathbf{k},\sigma}-\xi_{\mathbf{k}+\mathbf{q},\sigma}}$
$=-\frac{1}{N}\sum_{\mathbf{k}}\sum_{\sigma}\partial_{\epsilon}n_{\mathrm{FD}}(\xi_{\mathbf{k},\sigma}),$ (S17)
$\frac{1}{U}$ $=-\frac{1}{2}\sum_{\text{$\mathbf{k}$}}\sum_{\sigma}\lim_{q\rightarrow 0}\frac{n_{\mathrm{FD}}(\xi_{\mathbf{k},\sigma})-n_{\mathrm{FD}}(\xi_{\mathbf{k}+\mathbf{q},\bar{\sigma}})}{\xi_{\mathbf{k},\sigma}-\xi_{\mathbf{k}+\mathbf{q},\bar{\sigma}}}$
$=\frac{1}{N}\sum_{\mathbf{k}}\frac{n_{\mathrm{FD}}(\xi_{\mathbf{k},\uparrow})-n_{\mathrm{FD}}(\xi_{\mathbf{k},\downarrow})}{Um}.$ (S18)
Here $\partial_{\epsilon}n_{\mathrm{FD}}(\xi)\equiv\frac{\partial}{\partial\xi}n_{\mathrm{FD}}(\xi)$. The two equations determine the chemical potential and the magnetization, respectively.
The spin wave dispersion is determined from the pole of the spin susceptibility $\chi_{\mathrm{RPA}}^{ij}(\mathbf{q},\omega)$. We note that
$\chi_{0}^{+-}(\mathbf{q},\omega)=$ $-\frac{1}{2N}\sum_{\mathbf{k}}\sum_{\sigma}\frac{n_{\mathrm{FD}}(\xi_{\mathbf{k},\sigma})-n_{\mathrm{FD}}(\xi_{\mathbf{k}+\mathbf{q},\bar{\sigma}})}{\omega+\xi_{\mathbf{k},\sigma}-\xi_{\mathbf{k}+\mathbf{q},\bar{\sigma}}}$
$=$ $-\frac{1}{2N}\sum_{\mathbf{k}}\sum_{\sigma}\frac{n_{\mathrm{FD}}(\xi_{\mathbf{k},\sigma})-n_{\mathrm{FD}}(\xi_{\mathbf{k}+\mathbf{q},\bar{\sigma}})}{\omega+\epsilon_{\mathbf{k}}-\epsilon_{\mathbf{k}+\mathbf{q}}-Um}$
$\simeq$ $\frac{1}{U}+\frac{\omega}{mU^{2}}-\frac{q^{2}}{2m_{e}mU^{2}},$ (S19)
where we make the approximation $\epsilon_{\mathbf{k}}=\frac{\mathbf{k}^{2}}{2m_{e}}$ with $m_{e}$ being the effective mass of the electron. We obtain the spin susceptibility,
$\chi_{\mathrm{RPA}}^{ij}(\mathbf{q},\omega)=-\frac{\chi_{0}^{+-}(\mathbf{q},\omega)}{\frac{\omega}{mU}-\frac{q^{2}}{2m_{e}mU}}\delta_{ij}.$ (S20)
Therefore, the spin wave possesses the square dispersion $\omega(\mathbf{q})\propto q^{2}$.
V.2 Spin wave gap from magnetic anisotropy
Indeed, the introduction of magnetic anisotropy, which breaks the $SU(2)$ symmetry, can help stabilize the FM. A simple example of this is obtained by adding the anisotropic term (S5) to the Hubbard Hamiltonian. The $\delta H$ commutes with the Hubbard model Hamiltonian. Thus, the extra contribution to the Hartree-Fock Hamiltonian
$\delta H_{\mathrm{HF}}=-2zmJ_{3}\sum_{\mathbf{i}}S_{\mathbf{i}}^{z}+zm^{2}J_{3}$ (S21)
does not alter the Hartree-Fock ground state. Here, $z$ is the coordinate number. The spin susceptibility accordingly turns out to be
$\chi_{0}^{ij}(\mathbf{q},\omega)$ $\simeq\frac{1}{U}+\frac{\omega}{mU^{2}}-\frac{q^{2}/(2m_{e})+2zJ_{3}}{mU^{2}},$ (S22)
and
$\chi_{\mathrm{RPA}}^{ij}(\mathbf{q},\omega)=-\frac{\chi_{0}^{+-}(\mathbf{q},\omega)}{\frac{\omega}{mU}-\frac{q^{2}/(2m_{e})+2zJ_{3}}{mU}}\delta_{ij}.$ (S23)
which indicates a finite gap in Eq. (S6) of the spin wave. This gap arises from the breaking of $SU(2)$ symmetry due to the magnetic anisotropy. As a consequence, the finite gap $\Delta$ can suppress the thermal fluctuations of spin wave as such to stabilize a FM at finite temperature. The above RPA analysis can be applied to a general magnetic anisotropic term which can induce a spin-wave gap due to the breakdown of $SU(2)$ symmetry.